Suppose that a typical firm in a monopolistically competitive industry faces a demand curve given by:


q = 60 − (1/2)p, where q is quantity sold per week.


The firm’s marginal cost curve is given by: MC = 60.

  1. How much will the firm produce in the short run?
  2. What price will it charge?
1
Expert's answer
2021-10-04T10:19:45-0400

1. The profit-maximizing quantity produced in the short run is:

MR = MC,

p = 120 - 2q,

MR = TR'(q) = 120 - 4q,

120 - 4q = 60,

4q = 60,

q = 15 units.

2. The price it will charge is:

p = 120 - 2×15 = 90.